OpenStudy (anonymous):

A car factory produces 100 cars in a day. If c is the total number of cars produced in d days, which equation represents the number of cars as a function of the number of days? d = f(c) = 100c d = f(c) = 100 + c c = f(d) = 100d c = f(c) = 100d d = f(c) = 100 – d

OpenStudy (zale101):

If c is the \(\color{red}{total ~number}\) of cars produced in d days. What does this tell you? What is the output of the function, Is it c or d? What is the input of the function, is it c or d?

OpenStudy (zale101):

Member, a function is a set of rules that consists of an input and an output, for example y=f(x)=mx+b y is f(x) which is the output in the function. x is the input in the function.

OpenStudy (anonymous):

The output is d.

OpenStudy (anonymous):

Is it a ?

OpenStudy (anonymous):

@Zale101

OpenStudy (anonymous):

I already chose the 3rd one.

OpenStudy (zale101):

\(\color{red}y ~as~ a~ function ~of~ \color{blue}x\) is writing the equation as y=mx+b So, \(\color{red}{the~ number~ of~ cars}\) as a function of \color{blue}{ the~ number ~of ~days~\) y is the number of cars, and according to the question that represents c x is the number of days, and according to the question this represents d

OpenStudy (zale101):

c = f(d) = 100d ?

OpenStudy (zale101):

That's right

OpenStudy (anonymous):

Thank you! Can you help me out with another?

OpenStudy (zale101):

Have you tried working on it?

OpenStudy (anonymous):

Yes, actually.

OpenStudy (anonymous):

It just not adding up though. If your free and you don't mind helping, that'd be great!

OpenStudy (zale101):

Okay, then post a new problem.

OpenStudy (anonymous):

I'll tag you in it.

OpenStudy (anonymous):

Ok. i have tagged you!!